OpenStudy (anonymous):

find the slope intercept if x intercept is 3 and y is 4/5

OpenStudy (anonymous):

Are they only giving you that one point? If so, and you are assuming the line goes through the origin (0,0), then you cane find the slope by rise/run So... Slope = rise/run Slope = y/x

OpenStudy (anonymous):

ok it's not y = mx + b

OpenStudy (anonymous):

That is the overall equation, so m is your slope in that equation. and b is your y intercept. So if you know what b is then you could plug that in, otherwise if b goes to 0... y=mx+b assume b=0 y=mx m=y/x

OpenStudy (anonymous):

i thought b = 4/5

OpenStudy (anonymous):

Oh! So sorry, one second I misread the problem!

OpenStudy (anonymous):

You are correct! So b=4/5, y=mx+4/5 if you now use the other number, so when x=3, y=0 y=mx+4/5 0=3m+4/5 Now from there you can solve for m :) Sorry about the confusion